OpenStudy (anonymous):

MEDAL Which of the following statements are not true? Check all that apply. a. logM-logN=log(m-n) b. logM+logN=log(MN) c. logM^p=p*logM d. log(M/N)=logM/logN

OpenStudy (anonymous):

@ranga Do you know how to do this?

OpenStudy (shaik0124):

d option is not correct

OpenStudy (shaik0124):

correct answer for that is logM-logN

OpenStudy (shaik0124):

You have learned various rules for manipulating and simplifying expressions with exponents, such as the rule that says that x3 × x5 equals x8 because you can add the exponents. There are similar rules for logarithms. Log Rules: 1) logb(mn) = logb(m) + logb(n) 2) logb(m/n) = logb(m) – logb(n) 3) logb(mn) = n · logb(m) In less formal terms, the log rules might be expressed as: 1) Multiplication inside the log can be turned into addition outside the log, and vice versa. 2) Division inside the log can be turned into subtraction outside the log, and vice versa. 3) An exponent on everything inside a log can be moved out front as a multiplier, and vice versa. Warning: Just as when you're dealing with exponents, the above rules work only if the bases are the same. For instance, the expression "logd(m) + logb(n)" cannot be simplified, because the bases (the "d" and the "b") are not the same, just as x2 × y3 cannot be simplified (because the bases x and y are not the same).

OpenStudy (ranga):

a is wrong, b is correct Have you typed c correctly? There is no equal sign. d is wrong

OpenStudy (anonymous):

sry. I fixed it

OpenStudy (ranga):

c is correct. So a and d are wrong

OpenStudy (anonymous):

thank you.

OpenStudy (ranga):

yw.